Problem 10.10 Your answer is incorrect. Try again Six Twelve, Inc., is considering opening up a new convenience store in downtown New York City. The expected annual revenue at the new store is $680,000. To estimate the increase in working capital, analysts estimate the ratio of cash and cash-equivalents to revenue to be 0.03 and the ratios of receivables, inventories, and payables to revenue to be 0.05, 0.10, and 0.04, respectively, in the same industry. What is the expected incremental cash flow related to working capital when the store is opened? (Enter negative amount using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).) Incremental cash flow 95200
Answer #1

cash and cash equivalnet= .03*680000 = 20400

receivable = .05*680000 = 34000

inventories= .10*680000 = 68000

payables = .04*680000 = 27200

Total increase in current asset = 20400+34000+68000 = 122400

incremental cash flow = current asset -current liabilities

       = 122400-27200

       = 95200   [since it is a investment in working capital it should be entered as -95200]

incremental cash flow = -95200